OpenStudy (anonymous):

where do y=x^3 and y=1/x lines intersect ?

OpenStudy (anonymous):

1/x=x^3 1=x^4 solve x substitute x

OpenStudy (anonymous):

or you could always draw a graph

OpenStudy (anonymous):

thanks